OpenStudy (anonymous):

Which of the following is a solution to √x+10 +3=8? x = 15 x = 5 x = –5 x = –15

OpenStudy (johnweldon1993):

\[\large \sqrt{x} + 10 + 3 = 8\] like that?

OpenStudy (anonymous):

no, x+10 is under the radical sign, +3=8 is not

OpenStudy (johnweldon1993):

Ahh that makes sense... \[\large \sqrt{x + 10} + 3 = 8\] isolate the radical...by subtracting 3 from both sides... \[\large \sqrt{x + 10} = 5\] Now square both sides of the equation to get rid of the radical... \[\large x + 10 = 25 \] And finally subtract 10 from both sides of the equation... \[\large x = 15 \]

OpenStudy (johnweldon1993):

Hope that makes sense @skltngrl12 :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (johnweldon1993):

No problem :)